# Intuitive Fine-Tuning: Towards Unifying SFT and RLHF into a Single Process

### v1

@@ 3.4 From SFT to RLHF
$\forall n\in N:T_{\theta}(s_{n}^{*},\rho_{0})\leq T_{\theta}(s_{n}^{\theta},\rho_{0})\leq 1$ (18)
@@
secretly sets $T_{\theta}(s_{n-1}^{*},\rho_{0})=1$ during aligning $T_{\theta}(s_{n}^{*},\rho_{0})$ with $T^{*}(s_{n}^{*},\rho_{0})$. This makes an overestimation of the transition probabilities and preference of model, leading to an inferior optimization progress in SFT. Thus RLHF is needed for further preference alignment.
PPO enjoys an unbiased estimation of model preference, while employing a progressively unbiased estimation of human preference:
@@
Initially biased, this estimation gradually becomes unbiased as the model aligns with human preference over time.
As a reusult, PPO provides a closer approximation than SFT to the actual circumstances of model in Transition Optimization:
@@
which sets $T_{\theta}(s_{n}^{\theta},\rho_{0})=1$ and $\hat{s_{n}^{*}}=\pi^{*}(s_{n-1}^{\theta})$, at the expense of preference-labeling, reward modeling and online sampling for estimating $\pi^{*}(s_{n-1}^{\theta})$.
@@ 4 Method
While SFT is data and computation-efficient, it has an inferior approximation for both Preference Estimation and Transition Optimization. In the other side, RLHF (represented by PPO and DPO) enjoys better approximation at the expense of preference data construction. We hope to make good use of their strength, using solely the target data as SFT but having similar approximation as RLHF.
@@ 4.1 Intuitive Preference Estimation
A key distinction between SFT and RLHF lies in whether the full distribution of model preference for each initial instruction is sampled. Contrasted to RLHF, the intermediate state of the target answer used for prior in SFT may be far away from the model preference, leading to inferior outcomes.
@@ 4.2 Dynamic Relation Propagation
$\forall n\in[0,N]:\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (25)
$\hat{T_{\theta}}(s_{n}^{*},\rho_{0})$ $=\prod\limits_{i=0}^{n-1}T_{\theta}(s_{i+1}^{*},\hat{s_{i}^{\theta}})$ (26)
which can be optimized by the following loss function that quantifies the disparities of transition between model and human:
We make the same hypothesis as SFT that the optimization objective of each target intermediate state has a probability equal to 1, so that $\forall n\in[0,N]$:
$T^{*}(s_{n}^{*},\rho_{0})=1=T^{*}(s_{N}^{*},\rho_{0})$ (28)
@@
$\mathcal{L}(\hat{T_{\theta}}(s_{n}^{*},\rho_{0}))=-\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))$ (29)
which becomes easier to be implemented in parallel. See Appendix A.2 for complete derivation.
At the same time, we also prove that the objective optimized by this loss function implicitly satisfies the Bellman Equation in the on-policy case:
@@ 5 Experiments
We conduct experiments mainly on NLP tasks. Considering the absence of an optimal policy in NLP, we also utilize the Frozen Lake environment for further validation.
5.1 Settings
@@
Datasets. We select UltraChat-200k (Ding et al., 2023) and UltraFeedback-60k (Cui et al., 2023) as single-target and pair-wise dataset respectively.
@@
Models. We conduct experiments on Mistral-7B-v0.1 (Jiang et al., 2023) and its fine-tuned version Mistral-7B-sft-beta (Tunstall et al., 2023). The former serves as the base model, while the latter has been fine-tuned on the UltraChat-200k.
@@
Scenarios. We consider two different training scenarios, one is in RLHF-only mode and the other uses sequential recipe of SFT and RLHF. In the first scenario, alignment is conducted directly from base model Mistral-7B-v0.1 using UltraFeedback. In order to ensure balanced data volume between different method, we randomly sample 60k data from UltraChat as supplementary for SFT and IFT, for only the target data are utilized in these two methods. The second scenario is commonly seen, where SFT and RLHF is employed sequentially. For this scenario, we use Mistral-7B-sft-beta as start-point, which has been fine-tuned with UltraChat using SFT. Then we fine-tune it further with UltraFeedback using RLHF.
@@
Baselines. Our main baselines are SFT and DPO (Rafailov et al., 2024), with PPO excluded due to computational limitations. Additionally, we incorporate an improved versions of DPO, named ORPO (Hong et al., 2024), who claims to achieve alignment directly without SFT and reference model, aligning well with our goals. In addition to reproducing the algorithms mentioned above, we also consider Zephyr-7B-beta (Tunstall et al., 2023) and Mistral-ORPO-alpha (Hong et al., 2024), two open-source checkpoints that utilize sequential and direct recipes respectively. Both of them used start-point models and datasets similar to ours.
@@
Benchmarks. We consider two types of benchmarks. One is from the widely used Open-LLM LeaderBoard, which contains ARC-Challenge (25-shot) (Clark et al., 2018), MMLU (5-shot) (Chung et al., 2024), TruthfulQA (0-shot) (Lin et al., 2021), WinoGrande (5-shot) (Sakaguchi et al., 2021), and GSM8K (5-shot) (Cobbe et al., 2021). The other is LM-based evaluation, including Alpaca-Eval and Alpaca-Eval-2 (Dubois et al., 2024). We utilize chat template for all benchmarks to obtain a more accurate evaluation for chat models.
5.2 Main Results
@@
Effectiveness on Sequential Recipe. In this scenario, IFT demonstrates good performance across benchmarks having standard answers or not, see more details in table 2 and 4. On Open-LLM Leaderboard, IFT showcases the best average capabilities across all tasks, excelling particularly in tasks requiring generation, reasoning and fact-following abilities, such as TruthfulQA and GSM8K. However, IFT has a relatively large gap between DPO in multi-choice tasks like ARC-Challenge and MMLU. We will explained the reasons later. When evaluated for instruction-following and question-answering abilities as judged by GPT-4, IFT’s performance is comparable to that of the chosen baselines. Remarkably, IFT achieves these results using the least amount of data and computational resources among all the methods tested.
Method Alpaca-Eval Alpaca-Eval-2 win-rate lc win-rate win-rate lc win-rate Mistral-7B 24.72 11.57 1.25 0.35      with UltraFeedback-60k + SFT 82.56 78.32 7.09 8.67 + DPO 74.00 73.12 9.73 8.58 + ORPO 85.14 76.60 8.82 12.34 + IFT 85.18 78.78 9.95 13.27 Mistral-ORPO-$\alpha$ 87.92 11.33      with UltraChat-200k + UltraFeedback-60k sequentially + SFT 86.69 77.96 4.08 6.43 + SFT + SFT 86.34 76.98 4.55 7.14 + SFT + DPO 91.62 81.54 10.08 13.72 + SFT + ORPO 86.26 79.67 7.40 12.27 + SFT + IFT 88.37 81.29 10.26 14.34 Zephyr-7B-$\beta$ 90.60 10.99
@@
Method Data Volume GPU Memory Training Duration SFT v m t DPO 2v $\sim$2m $\sim$3t ORPO 2v $\sim$1.5m $\sim$2t IFT v m $\sim$1.5t
Table 3: Cost Estimation for Different Methods. We use the cost of SFT as the benchmark, where v, m, and t do not represent exact quantity units. We assume backward propagation needs higher cost than forward propagation in terms of GPU Memory and Training Duration.
@@
Table 4: Evaluation on Open-LLM Leaderboard with chat template. When fine-tuning with the same recipe, IFT achieves the highest average score across all methods. Directly conducting alignment using IFT showcases the best performance in all recipes with the least data and computation. In particular, IFT performs better in tasks requiring generation, reasoning and fact-following abilities, but falling behind in multi-choice tasks like ARC and MMLU.
Effectiveness on RLHF-only Recipe. IFT not only maintains the performance advantages compared with other baselines in this setting, as seen in the sequential scenario. But also, IFT performs comparably or even superiorly to many method in sequential recipe. While DPO tends to fail under this setting, ORPO remains competitive in its open-source model. However, when constrained in the same experiment setting, the performance of ORPO becomes worse than IFT. Additionally, the reliance on preference data makes ORPO more costly in terms of negative sampling, preference labeling, and GPU memory consumption. Consequently, IFT stands out as a more efficient and cost-effective alternative in this context.
Multi-Choice vs. Generation. When evaluated on Open-LLM Benchmark, IFT performs better on generation tasks but struggles with multi-choice tasks, whereas DPO exhibits the opposite performance. This discrepancy may be partially due to the different evaluation metrics between these two type of tasks, and different training objective between IFT and DPO. Multi-choice tasks are evaluated using the log-likelihood of model for each entire golden answer, while generation tasks require the model to build the final answer based on the selected token by itself ahead. So generation better reflects the causality and reasoning abilities than multi-choice. While DPO only focus on aligning the mapping between instructions and entire answers, IFT not only addresses this issue, but also places greater emphasis on the causal relationship between tokens. As a result, DPO tends to be expert on distribution mapping tasks like multi-choice, but IFT tends to perform better on tasks that requires model to explore token by token. Considering this, we changed ARC-Challenge into a generation task, during which the question and candidate answers would be displayed beforehand and answer would be extract from model’s generation. Without changing the distribution of the benchmark, IFT showcases advantage compared with any other method in this setting. Overall, IFT maintains the best balance across different tasks, achieving the highest average score.
Objective Trade-off between SFT and RLHF. Traditional RLHF methods deliver excellent alignment performance, particularly in enhancing the instruction-following ability of language models, as showed in table 2. However, fitting the different objectives of SFT and RLHF involves many hyper-parameter trade-offs. This phenomenon is also observed in table 4, where the models trained by sequential recipe of SFT and other RLHF methods showcase obvious inferior results on Open-LLM Leaderboard even worse than SFT alone. Avoiding this trade-off, ORPO and IFT can achieve better performance by directly conducting alignment on the base model.
@@
5.3 Analysis in Frozen-Lake Environment
Figure 3: The Frozen Lake Game. IFT performs much better than SFT and ORPO, slightly worse than DPO.
As scores on Open-LLM Leaderboard only partially reflect models’ performance, and GPT-4 inadequately models human language generation, further comparison to a truly optimal policy is necessary. Given the difficulty of obtaining an optimal policy representing human language, we validate our algorithm in a simplified setting called Frozen Lake (Farama, 2023). In this environment, an agent attempts to find a gift on a nearly frozen lake with several holes, terminating the game upon finding the gift or falling into a hole. The limited number of states and actions allows the optimal policy to be easily derived using classical RL methods.
@@ 5.3 Analysis in Frozen-Lake Environment
IFT achieves a significantly better policy than SFT and ORPO, although it performs slightly worse than DPO. This is partly because, in terms of comparing how closely the explored grid aligns with the agent’s preference, the order is DPO > IFT > ORPO > SFT. Although ORPO also considers the negative trajectories sampled from policy, its direct incorporation of SFT loss with a fusion coefficient deviates its preference estimation, partially diminishing its effectiveness. Additionally, DPO, ORPO and IFT explore more grids than SFT, which also helps the agent develop a better understanding of the environment.
@@ 6 Conclusion
In this paper, we first interpret SFT and some typical RLHF methods into a unified framework using Preference Estimation and Transition Optimization. We then introduce an efficient and effective method called Intuitive Fine-Tuning (IFT), which achieves alignment directly from the base model using non-preference-labeled data. Finally, experiments on widely used benchmarks in NLP setting and in Frozen Lake environment demonstrate the competitive performance of IFT.
Limitations and Future Work. Our validation of IFT is limited to the fine-tuning setting, where data volume is constrained, leaving the scalability of IFT unexplored. Additionally, we primarily use Mistral-7B for baseline testing, and the generalization of IFT to larger and more diverse models requires exploration in the future.
Acknowledgments
Special thanks to Yihao Liu, Che Jiang, and Xuekai Zhu for contributing to idea formulation, Jingkun Yang, and Xuanqi Dong for their assistance with mathematical definitions, and Hong Liu, and Chushu Zhou for insightful discussions during the experiments.
Appendix A Appendix

### v2

@@ 3.4 From SFT to Preference Optimization
$\forall n\in[0,N],s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*},s_{n}^{\theta}\in\mathcal{S}_{\rho_{0}}^{\theta}:T_{\theta}(s_{n}^{*},\rho_{0})\leq T_{\theta}(s_{n}^{\theta},\rho_{0})\leq 1$ (12)
@@
secretly sets $T_{\theta}(s_{n-1}^{*},\rho_{0})=1$ during aligning $T_{\theta}(s_{n}^{*},\rho_{0})$ with $T^{*}(s_{n}^{*},\rho_{0})$. This makes an overestimation of the transition probabilities and preference of model, leading to an inferior optimization progress in SFT. Thus Preference Optimization is needed for further preference alignment.
PPO shows an unbiased estimation of model preference, while employing a progressively unbiased estimation of human preference:
@@
Initially biased, this estimation gradually becomes unbiased as the model aligns with human preference over time. As $T_{\theta}(s_{n}^{\theta},\rho_{0})$ is consistently closer to 1 than $T_{\theta}(s_{n-1}^{*},\rho_{0})$, PPO provides an closer approximation than SFT to the actual circumstances of model in Transition Optimization:
@@
which sets $T_{\theta}(s_{n}^{\theta},\rho_{0})=1$ and $\hat{s_{n}^{*}}=\pi^{*}(s_{n-1}^{\theta})$. However, estimating $\pi^{*}(s_{n-1}^{\theta})$ is at the expense of preference-labeling, reward modeling and online sampling.
@@ 4 Method
While SFT is data and computation-efficient, it has an inferior approximation for both Preference Estimation and Transition Optimization. In the other side, Preference Optimization (represented by PPO and DPO) enjoys better approximation at the expense of preference data construction. We hope to make good use of their strength, using solely the target data as SFT but having similar approximation as Preference Optimization.
@@ 4.1 Intuitive Preference Estimation
A key distinction between SFT and Preference Optimization lies in whether the full distribution of model preference for each initial instruction is sampled. Contrasted to Preference Optimization, the intermediate state of the target answer used for prior in SFT may be far away from the model preference, leading to inferior outcomes.
@@ 4.2 Dynamic Relation Propagation
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:\hat{T_{\theta}}(s_{n}^{*},\rho_{0})\rightarrow T^{*}(s_{n}^{*},\rho_{0})$ (19)
where $s_{0}^{*}=\rho_{0}$ and $\hat{T_{\theta}}(s_{n}^{*},\rho_{0})=\prod\limits_{i=0}^{n-1}T_{\theta}(s_{i+1}^{*},\hat{s_{i}^{\theta}})$. This objective can be optimized by the following loss function that quantifies the disparities of transition between model and human:
$\mathcal{L}(\hat{T_{\theta}};T^{*})=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{n}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[\sum\limits_{n=0}^{N}\log\frac{\hat{T_{\theta}}(s_{n}^{*},\rho_{0})}{T^{*}(s_{n}^{*},\rho_{0})}\right]$ (20)
We make the same hypothesis as SFT that the optimization objective of each target intermediate state has a probability equal to 1:
$\forall n\in[0,N],\rho_{0}\in\mathcal{D},s_{n}^{*}\in\mathcal{S}_{\rho_{0}}^{*}:T^{*}(s_{n}^{*},\rho_{0})=1=T^{*}(s_{N}^{*},\rho_{0})$ (21)
@@
$\mathcal{L}_{\text{IFT}}(\mathcal{T}_{\theta},\delta_{\theta})=\mathbb{E}_{\rho_{0}\sim\mathcal{D}}\mathbb{E}_{s_{i}^{*}\sim\mathcal{S}_{\rho_{0}}^{*}}\left[-\sum\limits_{n=0}^{N}\sum\limits_{i=n}^{N}\log\mathcal{T}_{\theta}(a_{i}^{*},\delta_{\theta}(s_{i}^{*}))\right]$ (22)
where $a_{i}^{*}=\pi^{*}(s_{i}^{*})$. This reformulation facilitates parallel implementation, allowing IFT to achieve computational efficiency similar to SFT. See Appendix A.2 for complete derivation.
At the same time, we also prove that the objective optimized by this loss function implicitly satisfies the Bellman Equation for each state:
@@ 5 Experiments
We conduct experiments mainly on NLP setting. Considering the absence of an optimal policy of human language generation, we also utilize the Frozen Lake environment for further validation.
5.1 Settings for NLP
@@
Datasets. We select UltraChat-200k ding2023enhancing and UltraFeedback-60k cui2023ultrafeedback as single-target and pair-wise dataset respectively.
@@
Models. We conduct experiments on Mistral-7B-v0.1 jiang2023mistral and Mistral-7B-sft-beta tunstall2023zephyr , with the former as the base model and the latter fine-tuned on UltraChat-200k.
@@
Scenarios. We consider two different training scenarios, one using Preference Optimization exclusively, and the other employing sequential recipe of SFT and Preference Optimization. In the first scenario, alignment is conducted directly from base model Mistral-7B-v0.1 using UltraFeedback. In order to ensure balanced data volume between different method, we randomly sample 60k data from UltraChat as supplementary for SFT and IFT, for only the target data are utilized in these two methods. The second scenario is commonly seen, where SFT and Preference Optimization is employed sequentially. For this scenario, we use Mistral-7B-sft-beta as start-point, which has been fine-tuned with UltraChat using SFT. Then we fine-tune it further with UltraFeedback using Preference Optimization.
@@
Baselines. Our main baselines are SFT and DPO rafailov2024direct , with PPO excluded due to computational limitations. Additionally, we incorporate an improved versions of DPO, named ORPO hong2024reference , who claims to achieve alignment directly without SFT and reference model, aligning well with our goals. In addition to reproducing the algorithms mentioned above, we also consider Zephyr-7B-beta tunstall2023zephyr and Mistral-ORPO-alpha hong2024reference , two open-source checkpoints that utilize sequential and direct recipes respectively. Both of them used start-point models and datasets similar to ours.
@@
Benchmarks. We consider two types of benchmarks. One is from the widely used Open-LLM LeaderBoard, which contains ARC-Challenge (25-shot) clark2018think , MMLU (5-shot) chung2024scaling , TruthfulQA (0-shot) lin2021truthfulqa , WinoGrande (5-shot) sakaguchi2021winogrande , and GSM8K (5-shot) cobbe2021training . The other is LM-based evaluation, including Alpaca-Eval and Alpaca-Eval-2 dubois2024alpacafarm . We utilize chat template for all benchmarks to obtain a more accurate evaluation for chat models.
5.2 Main Results in NLP Tasks
@@
Effectiveness on Sequential Recipe. In this scenario, IFT demonstrates good performance across benchmarks having standard answers or not, see more details in Table 2 and 3. On Open-LLM Leaderboard, IFT showcases the best average capabilities across all tasks, excelling particularly in tasks requiring generation, reasoning and fact-following abilities, such as TruthfulQA and GSM8K. However, IFT has a relatively large gap between DPO in multi-choice tasks like ARC-Challenge and MMLU. We will explained the reasons later. When evaluated for instruction-following and question-answering abilities as judged by GPT-4, IFT’s performance is comparable to that of the chosen baselines. Remarkably, IFT achieves these results using the least amount of data and computational resources among all the methods tested.
@@
Method Reference Data Alpaca-Eval Alpaca-Eval-2 pairwise volume win-rate lc win-rate win-rate lc win-rate Mistral-7B 24.72 11.57 1.25 0.35      fine-tuning with UltraFeedback-60k + SFT 120k 82.56 78.32 7.09 8.67 + DPO 120k 74.00 73.12 9.73 8.58 + ORPO 120k 85.14 76.60 8.82 12.34 + IFT 120k 85.18 78.78 9.95 13.27 Mistral-ORPO-$\alpha$ 120k 87.92 11.33      fine-tuning with UltraChat-200k + UltraFeedback-60k sequentially + SFT 200k 86.69 77.96 4.08 6.43 + SFT + SFT 260k 86.34 76.98 4.55 7.14 + SFT + DPO 320k 91.62 81.54 10.08 13.72 + SFT + ORPO 320k 86.26 79.67 7.40 12.27 + SFT + IFT 260k 88.37 81.29 10.26 14.34 Zephyr-7B-$\beta$ 320k 90.60 10.99
Table 3: Evaluation on Open-LLM Leaderboard with chat template. When fine-tuning with the same recipe, IFT achieves the highest average score across all methods. Directly conducting alignment using IFT showcases the best performance in all recipes with the least data and computation.
@@
Effectiveness of Preference Optimization Alone. IFT not only maintains the performance advantages compared with other baselines in this setting, as seen in the sequential scenario. But also, IFT performs comparably or even superiorly to many method in sequential recipe. While DPO tends to fail under this setting, ORPO remains competitive in its open-source model. However, when constrained in the same experiment setting, the performance of ORPO becomes worse than IFT. Additionally, the reliance on preference data makes ORPO more costly in terms of negative sampling, preference labeling, and GPU memory consumption. Consequently, IFT stands out as a more efficient and cost-effective alternative in this context.
Multi-Choice vs. Generation. IFT performs better on generation tasks but struggles with multi-choice, whereas DPO exhibits the opposite performance. This may due to differences in evaluation metrics and training objectives zheng2023large ; plaut2024softmax ; tsvilodub2024predictions . Multi-choice tasks evaluate log-likelihood for entire answers, while generation tasks require token-by-token construction for causality and reasoning. DPO aligns the mapping between instructions and complete answers, while IFT emphasizes token-level causal relationships. As a result, DPO tends to excel in multi-choice tasks, while IFT performs better in token-by-token exploration tasks. In an ARC-Challenge adaptation to generation tasks, IFT demonstrates superiority without changing the benchmark’s distribution. Overall, IFT showcases its balanced performance across diverse tasks and achieving the highest average score.
Objective Trade-off between SFT and Preference Optimization. Traditional Preference Optimization methods deliver excellent alignment performance, particularly in enhancing the instruction-following ability of language models, as showed in Table 2. However, fitting the different objectives of SFT and Preference Optimization involves trade-offs tunstall2023zephyr . Even slight overfitting on SFT may result in reduced effectiveness of Preference Optimization. This phenomenon is also observed in Table 3, where the models trained by sequential recipe of SFT and other Preference Optimization methods showcase obvious inferior results on Open-LLM Leaderboard even worse than SFT alone. Avoiding this trade-off, ORPO and IFT can achieve better and more stable performance by directly conducting alignment on the base model.
@@
5.3 Further Validation in Frozen-Lake Environment
Figure 3: The Frozen Lake Game. Considering the MSE distance of parameters between the trained and optimal policy, IFT performs much better than SFT and ORPO, but slightly worse than DPO.
As scores on Open-LLM Leaderboard only partially reflect models’ performance, and GPT-4 inadequately models human language generation, further comparison to a truly optimal policy is necessary. Given the difficulty of obtaining an optimal policy representing human language, we validate our algorithm in a simplified setting called Frozen Lake frozenlake . In this environment, an agent attempts to find a gift on a nearly frozen lake with several holes, terminating the game upon finding the gift or falling into a hole. The limited number of states and actions in this game allows the optimal policy to be easily derived using classical RL methods.
@@ 5.3 Further Validation in Frozen-Lake Environment
In this setting, IFT achieves a significantly better policy than SFT and ORPO, although it performs slightly worse than DPO. This is partly because, in terms of comparing how closely the explored grid aligns with the agent’s preference, the order is DPO > IFT > ORPO > SFT. Although ORPO also considers the negative trajectories sampled from policy, its direct incorporation of SFT loss with a fusion coefficient deviates its preference estimation, partially diminishing its effectiveness. Additionally, DPO, ORPO and IFT explore more grids than SFT, which helps the agent develop a better understanding of the environment.
@@ 6 Conclusion
In this paper, we first interpret SFT and some typical Preference Optimization methods into a unified framework using Preference Estimation and Transition Optimization. We then introduce an efficient and effective method called Intuitive Fine-Tuning (IFT), which achieves alignment directly from the base model using non-preference-labeled data. Finally, experiments on widely used benchmarks in NLP setting and in Frozen Lake environment demonstrate the competitive performance of IFT.
Limitations and Future Work. Our validation of IFT is limited to the fine-tuning stage, where data volume is constrained, leaving the scalability of IFT unexplored. Additionally, we primarily use Mistral-7B for baseline testing, and the generalization of IFT to larger and more diverse models requires exploration in the future.
Acknowledgments and Disclosure of Funding
Special thanks to Yihao Liu, Che Jiang, and Xuekai Zhu for contributing to idea formulation, Pengfei Li, Xuanqi Dong, and Jingkun Yang for their assistance with mathematical definitions, Hong Liu, and Chushu Zhou for insightful discussions during the experiments.
Appendix A Theoretical Details